# Universal Physics Transformers: A Framework For Efficiently Scaling Neural Operators

### v2

@@ 4 Experiments
We ran experiments across different settings, assessing three key aspects of UPTs: (i) Effectiveness of the latent space representation. We test on steady state flow simulations in three dimensions, comparing against methods that use regular grid representations, and thus considerably larger latent space representations. (ii) Scalability. We test on transient flow simulations on large meshes. Specifically, we test the effectiveness of latent space rollouts, and assess how well UPTs generalize across different flow regime, and different domains, i.e., different number of mesh points and obstacles. (iii) Lagrangian dynamics modeling. Finally, we assess how well UPTs model underlying field characteristics when applied to particle-based simulations.
@@ 4.1 Steady state flows
[…]
In GINO, feature engineering in the form of a signed distance function (SDF) is used in addition to the mesh points to represent the irregular mesh as a regular grid. To include these features into UPT, we encode them into $5128^{3}=$ 512 latent SDF tokens using a shallow ConvNeXt V2 (Woo et al., 2023). These SDF tokens are concatenated to the latent tokens produced by the encoder and then fed to the approximator. When using the SDF features, we chose $n_{\text{latent}}=1024$ to balance the number of tokens from the mesh with the number of SDF tokens. Without SDF tokens, a much smaller number of tokens (64) is used. Implementation details and more information on the baseline models are provided in Appendix B.3.
ShapeNet-Car is a small-scale dataset. Consequently, methods that map the mesh onto a regular grid can employ grids of extremely high resolution, such that the number of grid points is orders of magnitude higher than the number of mesh points.. For example, a grid resolution of 64 points per spatial dimension results in 262.114 grid points, which is 73x the number of mesh points. As UPT is designed to operate directly on the mesh, we compare at different grid resolutions. The results in Tab. 2 demonstrate that UPT can model the underlying dynamics with a fraction of latent tokens and performs best across all grid sizes except at the highest resolution of $64^{3}$ where GINO performs slightly better.
Model SDF #Tokens MSE Mem. [GB]
U-Net 0 $64^{3}$ 6.13 1.3
FNO 0 $64^{3}$ 4.04 3.8
GINO 0 $64^{3}$ 2.34 19.8
UPT (ours) 0 $64$ 2.31 0.6
U-Net 32 $32^{3}$ 3.66 0.2
FNO 32 $32^{3}$ 3.31 0.5
GINO 32 $32^{3}$ 2.90 2.1
UPT (ours) 32 $10248^{3}+$ 1024 2.35 2.7
U-Net 64 $64^{3}$ 2.83 1.3
FNO 64 $64^{3}$ 3.26 3.8
GINO 64 $64^{3}$ 2.14 19.8
UPT (ours) 64 $10248^{3}+$ 1024 2.24 2.7
Table 2: Normalized test MSE for ShapeNet-Car pressure prediction. The loss is multiplied by 100. Memory denotes the amount required for a forward and backward pass of a single sample. UPTs can model the dynamics with a fraction of latent tokens compared to other models.
@@ 4.2 Transient flows
Figure 6: Transient flow results. MSE and correlation time on the testset. UPTs outperform compared methods on all model scales by a large margin.
Figure 7: Discretization convergence of the 68M parameter models from Fig. 6. Left: We vary the number of input/output points of models that were trained on inputs between 8K and 24K points and 8K target points. UPT demonstrates a stable performance across different number of input/outputs even if it has never seen that number of input/output points during training. Right: Increasing the number of supernodes improves the performance of UPT despite being trained with 2K supernodes. GINO was trained with 4K gridpoints.
@@
We test the scalability of UPTs on large-scale transient flow simulations. For this purpose, we self-generate 10K Navier-Stokes simulations within a pipe flow, which we split into 8K training, 1K validation and 1K test trajectories, using the pisoFoam solver from OpenFOAM (Weller et al., 1998), see Appendix A.3.2 for more information. For each simulation, between one and four objects (circles of variable size) are placed randomly within the pipe flow, and the uni-directional inflow velocity varies between 0.01 to 0.06 m/s. The temporal update of the numerical solver is initially set to $\Delta t=0.05$s. If instabilities occur, the trajectory is rerun with smaller $\Delta t$. Overall, each trajectory comprises 2K timesteps at the coarsest $\Delta t$ setting and 200K at the finest ($\Delta t=0.0005$s). After every second of simulated time, the corresponding timestep is written to the disk, resulting in 100 stored timesteps per trajectory. Each simulation contains between 29K and 59K mesh points where each point has three features: pressure, and the $x-$ and $y-$component of the velocity. A single simulation takes on average 120 seconds on 16 CPUs. In total, the dataset amounts to approximately 235GB of data in float16 precision.
@@
Model-wise, UPT uses the hierarchical encoder setup with all optional components depicted in Fig. 3. A message passing layer aggregates local information into $n_{\text{s}}=2048$ randomly selected supernodes, a transformer processes the supernodes and a perceiver pools the supernodes into $n_{\text{latent}}=512$ latent tokens. Approximator and decoder are unchanged. In Fig. 6, we compare UPT against GINO, U-Net and FNO. For U-Net and FNO, we interpolate the mesh onto a regular grid. We condition the models onto the current timestep and inflow velocity by modulating features within the model. We employ FiLM conditioning for U-Net (Perez et al., 2018), the “Spatial-Spectral” conditioning method introduced in Gupta & Brandstetter (2022) for FNO and GINO, and DiT for UPT (Peebles & Xie, 2023). Implementation details and more information on the baseline models are provided in Appendix B.4. We conduct a study to evaluate generalization to out-of-distribution object counts, velocities and geometry settings (e.g. replace circles with triangles) in Appendix B.6.
@@
We train all models for 100 epochs and evaluate test MSE as well as rollout performance for which we use the number of timesteps until the Pearson correlation of the rollout drops below 0.8 as evaluation metric (Kochkov et al., 2021). We do not employ any additional techniques to stabilize rollouts, see e.g., Lippe et al. (2024). Figure 6 shows that UPTs outperform compared methods on all model scales by a large margin. For all architectures, we didn’t observe much performance increase when further increasing parameter count, but rather observe that computation gets quickly infeasible with our current computational resources (UPT-68M and GINO-68M training takes roughly 450 A100 hours). The smaller variants UPT-8M and UPT-17M take roughly 150 and 200 A100 hours, respectively.
@@
While one would ideally use lots of supernodes and query the latent space with all positions during training, increasing those quantities increases training costs and the performance gains saturate. Therefore, we only use 2.048 supernodes and 16K randomly selected query positions during training. We investigate discretization convergence in the right part of Fig. 7 where we vary the number of input/output points and the number of supernodes. We use the 68M models without any retraining, i.e., we test models on “discretization convergence” as, during training, the mesh was discretized into 2.048 supernodes and 16K query positions. UPT generalizes across a wide range of different number of input or output positions, with even slight performance increases when using more input points. Similarly, using more supernodes increases performance slightly.
@@
Finally, we evaluate training with inverse encoding and decoding techniques, see Fig. 4. We investigate the impact of the latent rollout by training our largest model – a 68M UPT. The latent rollout achieves on par results to autoregressively unrolling via the physics domain, but speeds up the inference speed significantly as shown in Tab. 3. However, it is to note that in its current implementation the latent rollout requires a non-negligible overhead during training while greatly reducing costs and speed in inference.
@@ 4.3 Lagrangian fluid dynamics
Scaling particle-based methods such as discrete element methods or smoothed particle hydrodynamics to 10 million or more particles presents a significant challenge (Yang et al., 2020; Blais et al., 2019), yet it also opens a distinctive opportunity for neural surrogates. Such systems are far beyond the scope of this work. We however present a framing of how to model such systems via UPTs such that the studied scaling properties of UPTs could be exploited. In order to do so, we demonstrate how UPTs capture inherent field characteristics when applied to Lagrangian SPH simulations, as provided in LagrangeBench (Toshev et al., 2023b). Here, GNNs, such as Graph Network-based Simulators (GNS) (Sanchez-Gonzalez et al., 2020) and Steerable E(3) Equivariant Graph Neural Networks (SEGNNs) (Brandstetter et al., 2022b) are strong baselines, where predicted accelerations at the nodes are numerically integrated to model the time evolution of the particles. In contrast, UPTs learn underlying dynamics without dedicated particle-structures, and propagate dynamics forward without the guidance of numerical time integration schemes. An overview of the conceptual differences between GNS/SEGNN and UPTs is shown in Fig. 9.
@@
We use the Taylor-Green Vortex dataset in two and three dimensions (TGV2D, TGV3D). The Taylor-Green vortex (TGV) system was introduced by Taylor & Green (1937) as test scenario for turbulence modeling. The TGV system is an unsteady flow of a decaying vortex, displaying an exact closed form solution of the incompressible Navier–Stokes equations in Cartesian coordinates. We note that the TGV2D and TGV3D datasets model the same trajectory, however with different particles. Formulating the TGV system as UPT learning problem therefore means that the same trajectory is queried at different positions. Consequently, the evaluation against GNN-based simulators should be viewed as an illustration of the efficacy of UPTs in learning field characteristics, rather than a comprehensive GNN versus UPT comparison. For more information about the datasets see App. B.5.
@@
For UPT training, we input two consecutive velocities of the particles in the dataset at timesteps $t$ and $t-1$, and the respective particle positions. We regress two consecutive velocities at a later timesteps $\{t^{\prime}-1,t^{\prime}\}=\{t+\Delta T-1,t+\Delta T\}$ with mean-squared error (MSE) objective. For all experiments we use $\Delta T=10\Delta t$. The flexibility of the UPT encoder allows us to randomly sample $50\%$ up to $100\%$ of the total particles, diversifying the training procedure and forcing the model to learn the underlying dynamics. For both inverse encoding and decoding losses we use the velocities of all particles. We query the decoder to output velocities at target positions. UPTs encode the first two velocities of a trajectory, and autoregressively propagate dynamics forward in the latent space. We report the Euclidean norm of velocity differences across all $k$ particles. Figure 10 compares the rollout performance of GNS, SEGNN and UPT and shows the speedup of both methods compared to the SPH solver. The results demonstrate that UPTs effectively learn the underlying field dynamics while also managing much faster rollouts, with a 98-fold speedup compared to the SPH solver, a 55-fold speedup compared to SEGNN and a 11-fold speedup compared to GNS.
Model Time Speedup Device
SPH 4.9s 1x A6000
SEGNN 2.76s 2x A100
GNS 0.56s 9x A100
UPT 0.05s 98x A100
@@
Figure 10: (Top) Mean Euclidean norm of the velocity error over all particles for different timesteps. UPTs effectively learn the underlying field dynamics, resulting in lower velocity error as the trajectory evolves in time. (Bottom left) Visualization of the velocity field modeled by UPT (depicted in white) compared to the ground truth particle velocities. (Bottom right) Comparison of simulation/rollout runtimes for a TGV2D trajectory with 125 timesteps and 2500 particles across SPH simulation, GNS/SEGNN, and UPT, which has an inference time that results in a 98-fold speedup compared to SPH, a 55-fold speedup compared to SEGNN and a 11-fold speedup compared to GNS.
5 Future work
@@
We show that the latent rollout can be enabled via a simple end-to-end training procedure, but ultimately think that it can be improved by more delicate training procedures such as a two stage procedure akin to diffusion models (Rombach et al., 2022; Li et al., 2024). As there are many possible avenues to apply or improve the latent rollout, it would exceed the scope of this paper and we therefore leave it for future work.
@@
6 Conclusion
Acknowledgments
Appendix A Computational fluid dynamics
@@
A.1 Navier-Stokes equations
In our experiments, we mostly work with the incompressible Navier-Stokes equations (Temam, 2001). In two spatial dimensions, the evolution equation of the velocity flow field $\bm{u}:[0,T]\times U\subset\mathbb{R}^{3}\rightarrow\mathbb{R}^{2}$, $\bm{u}(t,x,y)=(u_{1}(t,x,y),u_{2}(t,x,y))$ with internal pressure $\bm{p}(x,y)=(p_{1}(x,y),p_{2}(x,y))$ and external force field $\bm{f}(x,y)=(f_{1}(x,y),f_{2}(x,y))$ is given via:
$\begin{split}\frac{\partial u_{1}}{\partial t}=&-u_{1}\frac{\partial u_{1}}{\partial x}-u_{2}\frac{\partial u_{1}}{\partial y}+\mu\left(\frac{\partial^{2}u_{1}}{\partial x^{2}}+\frac{\partial^{2}u_{1}}{\partial y^{2}}\right)-\frac{\partial p_{1}}{\partial x}+f_{1}\\ \frac{\partial u_{2}}{\partial t}=&-u_{1}\frac{\partial u_{2}}{\partial x}-u_{2}\frac{\partial u_{2}}{\partial y}+\mu\left(\frac{\partial^{2}u_{2}}{\partial x^{2}}+\frac{\partial^{2}u_{2}}{\partial y^{2}}\right)-\frac{\partial p_{2}}{\partial y}+f_{2}\\ 0=&\ \frac{\partial u_{1}}{\partial x}+\frac{\partial u_{2}}{\partial y}\end{split}$ (3)

### v3

@@ 4 Experiments
We ran experiments across different settings, assessing three key aspects of UPTs: (i) Effectiveness of the latent space representation. We test on steady state flow simulations in 3D, comparing against methods that use regular grid representations, and thus considerably larger latent space representations. (ii) Scalability. We test on transient flow simulations on large meshes. Specifically, we test the effectiveness of latent space rollouts, and assess how well UPTs generalize across different flow regime, and different domains, i.e., different number of mesh points and obstacles. (iii) Lagrangian dynamics modeling. Finally, we assess how well UPTs model the underlying field characteristics when applied to particle-based simulations. We outline the most important results in the following sections and provide implementation details and additional results in Appendix D. Most notabily, UPT also compares favorably against baseline regular grid methods on regular grid datasets D.3.
@@ 4.1 Steady state flows
For steady state prediction, we consider the ShapeNet-Car dataset generated by [105]. It consists of 889 car shapes from ShapeNet [20], where each car surface is represented by 3.6K mesh points in 3D space. We randomly create a train/test split containing 700/189 samples. We regress the pressure at each surface point with a mean-squared error (MSE) loss and sweep hyperparameters per model. Due to the small scale of this dataset, we train the largest possible model that is able to generalize the best. Training even larger models resulted in a performance decrease due to overfitting. We optimize the model size for all methods where the best mesh based models (GINO, UPT) contain around 300M parameters. The best regular grid based models (U-Net [86, 31], FNO [51]) are significantly smaller and range from 15M to 100M. Additional details are listed in Appendix D.4.
ShapeNet-Car is a small-scale dataset. Consequently, methods that map the mesh onto a regular grid can employ grids of extremely high resolution, such that the number of grid points is orders of magnitude higher than the number of mesh points. For example, a grid resolution of 64 points per spatial dimension results in 262K grid points, which is 73x the number of mesh points. As UPT is designed to operate directly on the mesh, we compare at different grid resolutions.
UPT is able to outperform models on smaller resolutions, for example UPT with only 64 latent tokens achieves a test MSE of 2.31 whereas GINO with resolution $48^{3}$ (110K tokens) achieves only 2.58 while taking significantly more runtime and memory. When additionally using feature engineering in the form of a signed distance function and $64^{3}$ grid resolution UPT achieves a competitive performance of 2.24 compared to 2.14 of GINO while remaining efficient. All test errors are multiplied by 100. UPT achieves a favorable cost-vs-performance tradeoff where the best GINO models requires 900 seconds per epoch whereas a only slightly worse (-0.17) UPT model takes only 4 seconds. We show a comprehensive table with all results in Appendix Tab. 4.
@@ 4.2 Transient flows
We test the scalability of UPTs on large-scale transient flow simulations. For this purpose, we self-generate 10K Navier-Stokes simulations within a pipe flow using the pisoFoam solver from OpenFOAM [110], which we split into 8K training, 1K validation and 1K test trajectories. For each simulation, between one and four objects (circles of variable size) are placed randomly within the pipe flow, and the uni-directional inflow velocity varies between 0.01 to 0.06 m/s. The simulation is carried out for 100s resulting in 100 timesteps that are used for training neural surrogates. Note that pisoFoam requires much smaller $\Delta t$ to remain stable (between 2K and 200K timesteps for 100s of simulation time). We use an adaptive meshing algorithm which results in 29K to 59K mesh points. Further dataset details are outlined in Appendix D.5.
@@
Model-wise, UPT uses the hierarchical encoder setup with all optional components depicted in Fig. 3. A message passing layer aggregates local information into $n_{\text{s}}=2048$ randomly selected supernodes, a transformer processes the supernodes and a perceiver pools the supernodes into $n_{\text{latent}}=512$ latent tokens. Approximator and decoder are unchanged. We compare UPT against GINO, U-Net and FNO. For U-Net and FNO, we interpolate the mesh onto a regular grid. We condition the models onto the current timestep and inflow velocity by modulating features within the model. We employ FiLM conditioning for U-Net [80], the “Spatial-Spectral” conditioning method introduced in [31] for FNO and GINO, and DiT for UPT [78]. Implementation details are provided in Appendix D.5.
@@
We train all models for 100 epochs and evaluate test MSE as well as rollout performance for which we use the number of timesteps until the Pearson correlation of the rollout drops below 0.8 as evaluation metric [43]. We do not employ any techniques to stabilize rollouts [57]. The left side of Fig. 5, shows that UPTs outperform compared methods by a large margin. Training even larger models becomes increasingly expensive and is infeasible for our current computational budget. UPT-68M and GINO-68M training takes roughly 450 A100 hours. UPT-8M and UPT-17M take roughly 150 and 200 A100 hours, respectively. Figure 4 shows a rollout and Appendix D.5.1 presents additional ones. We also study out-of-distribution generalization (e.g. more obstacles) in Appendix D.5.6.
Figure 5: Left and middle: MSE and correlation time on the testset. UPTs outperform compared methods on all model scales by a large margin. Right: We study discretization convergence by varying the number of input/output points or the number of gridpoints/supernodes of models that were trained on inputs between 8K and 24K points, 8K target points. UPT demonstrates a stable performance across different number of input/outputs even though it has never seen that number of input/output points during training. We study discretization convergence of supernodes in App. D.5.2, where UPT also shows a steady improvement when more supernodes are used during inference. Additionally, we study smaller UPT models and training on less data in Appendix D.5.4 and D.5.5.
@@
While one would ideally use lots of supernodes and query the latent space with all positions during training, increasing those quantities increases training costs and the performance gains saturate. Therefore, we only use 2048 supernodes and 16K randomly selected query positions during training. We investigate discretization convergence in the right part of Fig. 5 where we vary the number of input/output points and the number of supernodes. We use the 68M models without any retraining, i.e., we test models on “discretization convergence” as, during training, the mesh was discretized into 2048 supernodes and 16K query positions. UPT generalizes across a wide range of different number of input or output positions, with even slight performance increases when using more input points. Similarly, using more supernodes increases performance slightly. Additionally, we investigate training with more supernodes and/or more latent tokens in a reduced setting in Appendix D.5.3.
@@
Finally, we evaluate training with inverse encoding and decoding techniques, see Fig. 3. We investigate the impact of the latent rollout by training our largest model – a 68M UPT. The latent rollout achieves on par results to autoregressively unrolling via the physics domain, but speeds up the inference significantly as shown in Tab. 1. However, in its current implementation the latent rollout requires a non-negligible overhead during training. We discuss this limitation in Appendix A.
@@ 4.3 Lagrangian fluid dynamics
Scaling particle-based methods such as discrete element methods or smoothed particle hydrodynamics to 10 million or more particles presents a significant challenge [117, 9], yet it also opens a distinctive opportunity for neural surrogates. Such systems are far beyond the scope of this work. We however present a framing of how to model such systems via UPTs such that the studied scaling properties of UPTs could be exploited. In order to do so, we demonstrate how UPTs capture inherent field characteristics when applied to Lagrangian SPH simulations, as provided in LagrangeBench [101]. Here, GNNs, such as Graph Network-based Simulators (GNS) [87] and Steerable E(3) Equivariant Graph Neural Networks (SEGNNs) [14] are strong baselines, where predicted accelerations at the nodes are numerically integrated to model the time evolution of the particles. In contrast, UPTs learn underlying dynamics without dedicated particle-structures, and propagate dynamics forward without the guidance of numerical time integration schemes. An overview of the conceptual differences between GNS/SEGNN and UPTs is shown in Fig. 6.
@@
We use the Taylor-Green vortex dataset in three dimensions (TGV3D). The TGV system was introduced as a test scenario for turbulence modeling [97]. It is an unsteady flow of a decaying vortex, displaying an exact closed form solution of the incompressible Navier–Stokes equations in Cartesian coordinates. We note that the TGV3D dataset models the same trajectories but does so by tracking particles using SPH to solve the equations. Formulating the TGV system as a UPT learning problem allows the same trajectories to be queried at different positions, enabling the recovery of the complete velocity field, whereas GNNs can only evaluate the velocities of the particles. Consequently, the evaluation against GNNs should be viewed as an illustration of the efficacy of UPTs in learning field characteristics, rather than a comprehensive GNN versus UPT comparison. More details and experiments on the 2D version of the Taylor-Green vortex dataset (TGV2D) are in Appendix D.6.
@@
For UPT training, we input two consecutive velocities of the particles in the dataset at timesteps $t$ and $t-1$, and the respective particle positions. We regress two consecutive velocities at a later timesteps $\{t^{\prime}-1,t^{\prime}\}=\{t+\Delta T-1,t+\Delta T\}$ with mean-squared error (MSE) objective. For all experiments we use $\Delta T=10\Delta t$. We query the decoder to output velocities at target positions. UPTs encode the first two velocities of a trajectory, and autoregressively propagate dynamics forward in the latent space. We report the Euclidean norm of velocity differences across all $k$ particles. Figure 7 compares the rollout performance of GNS, SEGNN and UPT and shows the speedup of both methods compared to the SPH solver. The results demonstrate that UPTs effectively learn the underlying field dynamics.
@@
Figure 7: Left: Velocity error over all particles for different timesteps. UPTs effectively learn the underlying field dynamics, resulting in lower error as the trajectory evolves in time. Right: Visualization of the velocity field modeled by UPT (white) vs the ground truth particle velocities.
5 Discussion
Potential for extreme scale
UPTs consist of mainly transformer [106] layers which allows application of the same scaling and parallelism principles as are commonly used for training extreme-scale language models. For example, the recent work Llama 3 [27] trained on up to 16K H100 GPUs. While we do not envision that we train UPT on such a massive scale in the foreseeable future, the used techniques can be readily applied to UPTs.
Benefits of the latent rollout
While the latent rollout does not provide a significant performance improvement, it is almost an order of magnitude faster. The UPT framework allows to trade-off training compute vs inference compute. If inference time is crucial for a given application, one can train UPT with the inverse encoding and decoding objectives, requiring more training compute but greatly speeding up inference. If inference time is not important, one can simply train UPT without the reconstruction objectives to reduce training costs.
Additionally, the latent rollout enables applicability to Lagrangian simulations. As UPT models the underlying field instead of tracking individual particle positions it does not have access to particle locations at inference time. Therefore, autoregressive rollouts are impossible since the encoder requires particle positions as input. When using the latent rollout, it is sufficient to know the initial particle positions as dynamics are propagated without any spatial positions. After propagating the latent space forward in time, the latent space can be queried at arbitrary positions to evaluate the underlying field at given positions. We show this by querying with regular grid positions in Figure 7.
6 Conclusion
Acknowledgments
Appendix A Limitations and Future work
@@
We show that the latent rollout can be enabled via a simple end-to-end training procedure, but ultimately think that it can be improved by more delicate training procedures such as a two stage procedure akin to diffusion models [85, 55]. As there are many possible avenues to apply or improve the latent rollout, it would exceed the scope of this paper and we therefore leave it for future work.
@@
Appendix B Computational fluid dynamics
@@
B.1 Navier-Stokes equations
In our experiments, we mostly work with the incompressible Navier-Stokes equations [98]. In two spatial dimensions, the evolution equation of the velocity flow field $\bm{u}:[0,T]\times U\subset\mathbb{R}^{3}\rightarrow\mathbb{R}^{2}$, $\bm{u}(t,x_{1},x_{2})=(u_{1}(t,x_{1},x_{2}),u_{2}(t,x_{1},x_{2}))$ with internal pressure $\bm{p}(x_{1},x_{2})=(p_{1}(x_{1},x_{2}),p_{2}(x_{1},x_{2}))$ and external force field $\bm{f}(x_{1},x_{2})=(f_{1}(x_{1},x_{2}),f_{2}(x_{1},x_{2}))$ is given via:
$\begin{split}\frac{\partial u_{1}}{\partial t}=&-u_{1}\frac{\partial u_{1}}{\partial x_{1}}-u_{2}\frac{\partial u_{1}}{\partial x_{2}}+\mu\left(\frac{\partial^{2}u_{1}}{\partial x_{1}^{2}}+\frac{\partial^{2}u_{1}}{\partial x_{2}^{2}}\right)-\frac{\partial p_{1}}{\partial x_{1}}+f_{1}\\ \frac{\partial u_{2}}{\partial t}=&-u_{1}\frac{\partial u_{2}}{\partial x_{1}}-u_{2}\frac{\partial u_{2}}{\partial x_{2}}+\mu\left(\frac{\partial^{2}u_{2}}{\partial x_{1}^{2}}+\frac{\partial^{2}u_{2}}{\partial x_{2}^{2}}\right)-\frac{\partial p_{2}}{\partial x_{2}}+f_{2}\\ 0=&\ \frac{\partial u_{1}}{\partial x_{1}}+\frac{\partial u_{2}}{\partial x_{2}}\end{split}$ (3)